OpenStudy (anonymous):

Write the comparison as a ratio in simplest form, using a fraction. 0.77 to 0.35

OpenStudy (anonymous):

0.77/0.35 = 2 therefore 2, being the factor of 0.35 to 0.77, can be explained with the fraction 2/1 or 2:1 as a ratio

OpenStudy (anonymous):

how do you solve these type of problems

OpenStudy (anonymous):

with division or multiplication, and other fractions you also have to realise that whole numbers can be the same as a percentage * 100 so 1 would be 100% and 2 is 200% etc so if you know 70/35 = 2 then you can work out 77/35 in a ratio sorry about this explanation, it means that 0.77 to 0.35 isnt 2:1... but you might be able to work out what it actually is with that explanation i just gave. also you have to understand what multiplication and division actually are

OpenStudy (anonymous):

this made no sense to me lol

OpenStudy (anonymous):

its actually 11:5 in case you can't do it i just did 0.77/0.35 which = 2.2 so multiply 2.2*5 then you get 11 and multiply 1*5 to get 5 so your ratio is now 11:5

OpenStudy (anonymous):

if i was to give you a general way to figure these out it would be: take the first number, divide it by the second... that gives the difference between the first and second number... (you can multiply the number you just calculated by the 2nd number you used to get the first) this means that the second number is "so many times bigger or smaller than the first number used"

OpenStudy (anonymous):

sorry about my poor explanationsss but i think if you have the time, you can work your way through this: http://www.mathcentre.ac.uk/resources/workbooks/mathcentre/ratios.pdf and you'll be pro at it by the end (really good)

OpenStudy (anonymous):

see we have 0.77 to 0.35 now if you learnt we can write 0.77 as 77/100 and 0.35 as 35/100 0.77:0.35= 77/100/35/100 = 77/35 ans